Β If the probability mass function of a random variable is given by 𝑃(𝑋 = π‘Ÿ) = π‘˜π‘Ÿ 3 , π‘Ÿ = 1,2,3, 𝑓, 𝑓𝑖𝑛𝑑 (i)the value of k, (ii) 𝑃( 1 2 < 𝑋 < 5 2 /𝑋 > 1 3 ), (iii) the mean and variance of X.


1
Expert's answer
2021-04-14T10:50:22-0400

Given function,

"P(X=r)=kr^3, r=1,2,3" .


The probability Distribution is given as-





(i) As we know sum of probabilitied=1

"\Rightarrow k+8k+27k=1\Rightarrow 36k=1\Rightarrow k=0.028"


(ii)"P( \dfrac{(1.2<X<5.2)}{X>1.3})=\dfrac{0.224+0.756}{0.224+0.756}=\dfrac{0.98}{0.98}=1"


(iii) Mean "=\sum XP(X)=2.744"

Variance "=\sum X^2P(X)=7.728"